OpenStudy (anonymous):

A ball is dropped from the top of a tower 100m high simultaneously another ball is thorwn upward with a speed of 50 meter per second. After what time do they cross each other? (a) 1s (b) 2s (c) 3s (4) 4s

OpenStudy (anonymous):

When they cross each other they will have the same displacement. Set up equations for the displacements of each and equate them.

OpenStudy (anonymous):

but which equation, i've tried by appling S=vi t+1/2 at^2 but that does not work for me.Please can you solve it?

OpenStudy (anonymous):

You are on the right track. For the ball dropped downward, vi = 0. So\[\color{green}{\vec{d} = \frac{1}{2}\vec{g}t^2 }\] Set up the other equation and equate the two and solve for t.

OpenStudy (anonymous):

i've tried many methods but couldn't solve it,please solve it if you can

OpenStudy (pawanyadav):

Assume displacemt x and 100-x for thrown and release one. Then 100-x=0.5gt^2 x=100-5t^2 Again For the thrown ball x=50t-5t^2 As time for both balls are equal Equating both values you will get the answer 2.

OpenStudy (shamim):

Did u get it?

OpenStudy (shamim):

We wanna help u. Response plz!

OpenStudy (anonymous):

ok,thanks i've got it